OpenStudy (anonymous):

Can someone explain to me how to solve for x if i have 1/3 + 1/2 = 1/x

OpenStudy (anonymous):

determine what \[\frac{ 1 }{ 3} + \frac{ 1 }{ 2 } \] equals to first

OpenStudy (anonymous):

You can do this by changing the denominator to a common multipy of 2 and 3.

OpenStudy (anonymous):

so it would be 6/5?

OpenStudy (anonymous):

Are you supposed to add across the bottom?

OpenStudy (anonymous):

@Raunak401

OpenStudy (anonymous):

wait no.. \[\frac{ 1 }{ 3 }+\frac{ 1 }{ 2 }=\frac{ 5 }{ 6 } \] so what do i do from there?

OpenStudy (anonymous):

@ganeshie8